OpenStudy (anonymous):

Choose numbers a and b that make the function f (x) = { 3/4 0 a valid continuous probability distribution. ...if x is between a and b everywhere else

OpenStudy (anonymous):



OpenStudy (anonymous):

The function is very unclear... is it \[f(x)=\begin{cases}\dfrac{3}{4}&\text{for }a\le x\le b\\\\ 0&\text{elsewhere}\end{cases}~~~~?\] If it is, you must satisfy the following integral: \[\int_a^bf(x)~dx=\int_a^b\frac{3}{4}~dx=1\] Integrating, you have \[\frac{3}{4}(b-a)=1\] Now, since you have two unknowns, you must choose one as a starting point. The simplest might be \(a=0\), so that \[\frac{3}{4}(b-0)=1~~\iff~~b=\frac{4}{3}\] So, \[f(x)=\begin{cases}\dfrac{3}{4}&\text{for }0\le x\le \dfrac{4}{3}\\\\ 0&\text{elsewhere}\end{cases}\] Of course, that's not the only possible answer. You could have picked \(a=1\), which would make \(b=\dfrac{7}{3}\). In general, if \(a=n\), then \(b=\dfrac{4}{3}+n\).

OpenStudy (anonymous):

Thank you very much. I will work with what you have given me!

OpenStudy (anonymous):

yw